OpenStudy (anonymous):

HELPP! Solve -8(y+4)=10y+4

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Hehee?

OpenStudy (anonymous):

luis has his ways ^_-

OpenStudy (anonymous):

Ahhh

OpenStudy (anonymous):

-8(y+4)=10y+4 -8y-32=10y+4 -32-4=10y+8y

Directrix (directrix):

-8(y+4)=10y+4 -8y -32 = 10y + 4 -8y - 10y = 32 + 4 -18y = 36 y = -2

OpenStudy (anonymous):

Huh? I got y=.2..

OpenStudy (anonymous):

it's -2

OpenStudy (anonymous):

....